# Statistical Inference in Reinforcement Learning: A Selective Survey

### v1

@@ 4 Methodology: Forward-backward (Generative) Learning
    • Chen and Hong (2012) used local polynomial regression to estimate the conditional expectation. They employed undersmoothing – which reduces the bias of the test statistic at the cost of increased variance – to theoretically control the type-I error. However, in practice, their test often fails to achieve consistency in moderate to high-dimensional settings; see the simulation results in Zhou et al. (2023) for details.
  • Double-regression-based tests: Zhang et al. (2019), Shah and Peters (2020) and Quinzan et al. (2023) proposed double-regression-based tests for conditional independence. Their test statistics are doubly robust as well. While conceptually similar to Shi et al. (2020) and Zhou et al. (2023), these methods focus solely on conditional independence instead of the Markov assumption. Moreover, their test functions $g$ and $h$ are limited to identity functions, which simplifies the testing but reduces its power; see Shi et al. (2021) and Shi et al. (2023) for empirical comparisons with these double-regression-based tests in the context of conditional independence testing.
  • Strong conditional independence tests: The proposed forward-backward-learning-based tests primarily focus on weak conditional independence (Daudin, 1980) and cannot detect all alternative hypotheses. In other words, there exist alternative hypotheses under which the left-hand-side of Equation (1) is 0 for all $q$, $g$, and $h$. To address this, strong conditional independence tests can be developed to detect all alternatives; see the discussion following Theorem 5 in Shi et al. (2021). However, while weak conditional independence tests may fail to detect certain alternatives, they often have larger power in most practical scenarios.
@@ 5 Off-policy Confidence Interval Estimation
Off-policy evaluation (OPE) aims to assess the impact of implementing a newly developed target policy $\pi$ based on a historical dataset collected from a possibly different behavior policy. There is an extensive literature on OPE; see Uehara et al. (2022) for a recent review. OPE is particularly relevant to statistical inference, as in many applications, we are not only interested in the point estimator of the target policy $\pi$’s expected return $J(\pi)$, but also require uncertainty quantification for this estimator. Existing methodologies can be roughly divided into the following categories, depending on the type of point estimator used to estimate $J(\pi)$:
  1. 1.
    Model-based method: Hanna et al. (2017) adopted a model-based approach to derive the estimated $J(\pi)$ by learning the underlying MDP model and applied the bootstrap method to construct this value estimator’s associated confidence interval (CI);
  2. 2.
    Direct method: Feng et al. (2020), Luckett et al. (2020), Hao et al. (2021), Liao et al. (2021) and Shi et al. (2022) employed the direct method to learn $J(\pi)$ based on an estimated value or Q-function, and used concentration inequalities, asymptotic approximations or bootstrap to derive the CI based on their direct method estimator;
  3. 3.
    Important sampling (IS): Thomas et al. (2015) and Dai et al. (2020) studied IS-type estimators that reweight the immediate reward using IS ratios to address the distributional shift between target policy and behavior policy. They proposed the associated CIs based on concentration inequalities and the empirical likelihood approach (Owen, 2001), respectively;
  4. 4.
@@
    Doubly robust method (DR): DR combines the direct method and IS for more robust estimation of $J(\pi)$. A number of DR estimators have been proposed in the literature (see e.g., Zhang et al., 2013; Jiang and Li, 2016; Thomas and Brunskill, 2016; Uehara et al., 2020; Kallus and Uehara, 2022; Liao et al., 2022), and CIs can be readily constructed based on these estimators. Specifically, under the Markov assumption, when the estimated IS ratio and Q-function converge at a rate faster than $O_{p}(N^{-1/4})$, these estimators are asymptotically equivalent to sums of martingale differences, which behave similarly to i.i.d. sums. As a result, the standard sampling variance estimator can be used to quantify their standard error, enabling the resulting Wald-type CI to achieve nominal coverage. When the IS ratio or Q-estimator fails to achieve this convergence rate, high-order influence functions can be employed to further reduce the bias of the resulting estimator and ensure the validity of the Wald-type CI (Shi et al., 2021).
@@
To conclude this section, we remark that nearly all the aforementioned work assumed that the target policy $\pi$ is known in advance. In settings where $\pi$ corresponds to the optimal policy $\pi^{*}$, which needs to be estimated, constructing confidence intervals becomes particularly challenging in nonregular cases where $\pi^{*}$ is non-unique. Several methods have been proposed in the statistics literature to address this issue, including the $m$-out-of-$n$ bootstrap (Chakraborty et al., 2014), the online one-step method (Luedtke and Van Der Laan, 2016), and subagging (Shi et al., 2020).
@@ 6 Discussion
  • Testing the stationarity assumption: This paper primarily focuses on hypothesis testing of the Markov assumption in RL, assuming stationarity of the observation-action-reward triplets over time. However, the stationarity assumption itself can be violated. In the literature, several studies have proposed tests for the stationarity assumption, covering both model-free (Padakandla et al., 2020; Li et al., 2025) and model-based approaches (Alegre et al., 2021; Wang et al., 2023). Results from these studies demonstrate that such tests enable accurate change point detection in nonstationary environments, ultimately enhancing the performance of the subsequent policy optimization.
  • A/B testing: A/B testing is widely used in technological industries for comparing the performance of a newly developed policy against a standard control during product deployment (see Larsen et al., 2024; Quin et al., 2024, for reviews). It is inherently a statistical inference problem that computes $p$-values to assess the significance of improvements achieved by the new policy relative to the standard one. Recently, there has been growing interest in integrating RL into A/B testing to evaluate long-term treatment effects, which are common in practice (Glynn et al., 2020; Farias et al., 2022; Li et al., 2023; Shi et al., 2023; Wen et al., 2024). This emerging research direction warrants further exploration.

### v2

@@ 4 Methodology: Forward-backward (Generative) Learning
    • Chen and Hong (2012) used local polynomial regression to estimate the conditional expectation. They employed undersmoothing – which reduces the bias of the test statistic at the cost of increased variance – to theoretically control the type-I error. However, in practice, their test often fails to achieve consistency in moderate to high-dimensional settings; see the simulation results in Zhou et al. (2023b) for details.
  • Double-regression-based tests: Zhang et al. (2019), Shah and Peters (2020) and Quinzan et al. (2023) proposed double-regression-based tests for conditional independence. Their test statistics are doubly robust as well. While conceptually similar to Shi et al. (2020b) and Zhou et al. (2023b), these methods focus solely on conditional independence instead of the Markov assumption. Moreover, their test functions $g$ and $h$ are limited to identity functions, which simplifies the testing but reduces its power; see Shi et al. (2021b) and Shi et al. (2023b) for empirical comparisons with these double-regression-based tests in the context of conditional independence testing.
  • Strong conditional independence tests: The proposed forward-backward-learning-based tests primarily focus on weak conditional independence (Daudin, 1980) and cannot detect all alternative hypotheses. In other words, there exist alternative hypotheses under which the left-hand-side of Equation (1) is 0 for all $q$, $g$, and $h$. To address this, strong conditional independence tests can be developed to detect all alternatives; see the discussion following Theorem 5 in Shi et al. (2021b). However, while weak conditional independence tests may fail to detect certain alternatives, they often have larger power in most practical scenarios.
@@
Table 4.2: A summary of existing methodologies for off-policy confidence interval estimation.
Model-based Direct method Importance sampling Doubly robust
Concentration Feng et al. (2020) Thomas et al. (2015) Thomas and Brunskill (2016)
inequalities Jiang and Li (2016)
Zhou et al. (2023a)
Normal Luckett et al. (2020) Wang et al. (2023a) Shi et al. (2021a)
approximation Liao et al. (2021) Liao et al. (2022)
Shi et al. (2022) Kallus and Uehara (2022)
Bootstrap Hanna et al. (2017) Hao et al. (2021) Thomas and Brunskill (2016)
Hanna et al. (2017)
Empirical Dai et al. (2020)
likelihood
@@ 5 Off-policy Confidence Interval Estimation
Off-policy evaluation (OPE) aims to assess the impact of implementing a newly developed target policy $\pi$ based on a historical dataset collected from a possibly different behavior policy. There is an extensive literature on OPE; see Uehara et al. (2022) for a recent review. OPE is particularly relevant to statistical inference, as in many applications, we are not only interested in the point estimator of the target policy $\pi$’s expected return $J(\pi)$, but also require uncertainty quantification for this estimator.
We summarize the existing OPE methodologies in Table 4.2, which can be broadly categorized into four groups based on the approach used to construct confidence intervals (CIs). Each group corresponds to a row in the table and reflects a specific type of methods for CI construction: concentration inequalities (Zhang and Chen, 2020), normal approximation (Casella and Berger, 2024), bootstrap (Efron and Tibshirani, 1994), and empirical likelihood (Owen, 2001).
Meanwhile, these methodologies can also be classified according to the type of point estimator used to estimate $J(\pi)$, as summarized across the columns of the table. depending on the type of point estimator used to estimate $J(\pi)$. Specifically, model-based method derives the estimated $J(\pi)$ by learning the underlying MDP model; the direct method learns $J(\pi)$ based on an estimated value or Q-function; Importance sampling (IS) reweight the immediate reward using IS ratios to address the distributional shift between target policy and behavior policy; and doubly robust (DR) method combines the direct method and IS for more robust estimation of $J(\pi)$.
@@
We use DR (see e.g., Zhang et al., 2013; Jiang and Li, 2016; Thomas and Brunskill, 2016; Uehara et al., 2020; Kallus and Uehara, 2022; Liao et al., 2022) as an example to elaborate on the CI construction. Under the Markov assumption, when the estimated IS ratio and Q-function converge at a rate faster than $O_{p}(N^{-1/4})$, the DR estimator is asymptotically equivalent to a sum of martingale differences, which behaves similarly to i.i.d. sums. As a result, the standard sampling variance estimator can be used to quantify their standard error, enabling the resulting Wald-type CI to achieve nominal coverage. When the IS ratio or Q-estimator fails to achieve this convergence rate, high-order influence functions can be employed to further reduce the bias of the resulting estimator and ensure the validity of the Wald-type CI (Shi et al., 2021a).
@@
To conclude this section, we remark that nearly all the aforementioned work assumed that the target policy $\pi$ is known in advance. In settings where $\pi$ corresponds to the optimal policy $\pi^{*}$, which needs to be estimated, constructing confidence intervals becomes particularly challenging in nonregular cases where $\pi^{*}$ is non-unique. Several methods have been proposed in the statistics literature to address this issue, including the $m$-out-of-$n$ bootstrap (Chakraborty et al., 2014), the online one-step method (Luedtke and Van Der Laan, 2016), and subagging (Shi et al., 2020a).
@@ 6 Discussion
  • Testing the stationarity assumption: This paper primarily focuses on hypothesis testing of the Markov assumption in RL, assuming stationarity of the observation-action-reward triplets over time. However, the stationarity assumption itself can be violated. In the literature, several studies have proposed tests for the stationarity assumption, covering both model-free (Padakandla et al., 2020; Li et al., 2025) and model-based approaches (Alegre et al., 2021; Wang et al., 2023b). Results from these studies demonstrate that such tests enable accurate change point detection in nonstationary environments, ultimately enhancing the performance of the subsequent policy optimization.
  • A/B testing: A/B testing is widely used in technological industries for comparing the performance of a newly developed policy against a standard control during product deployment (see Larsen et al., 2024; Quin et al., 2024, for reviews). It is inherently a statistical inference problem that computes $p$-values to assess the significance of improvements achieved by the new policy relative to the standard one. Recently, there has been growing interest in integrating RL into A/B testing to evaluate long-term treatment effects, which are common in practice (Glynn et al., 2020; Farias et al., 2022; Li et al., 2023a; Shi et al., 2023a; Wen et al., 2024). This emerging research direction warrants further exploration.
  • Statistical inference in bandits: Finally, bandit is a special class of RL problems, characterized by independent state transitions. There has been a growing literature focused on statistical inference for adaptively collected data using bandit algorithms (Zhang et al., 2020; Bibaut et al., 2021; Dimakopoulou et al., 2021; Hadad et al., 2021; Simchi-Levi and Wang, 2023).